OpenStudy (january123):

1. The dot plot below shows how many apple slices each girl at Kayla's party ate. What is the mean of the data set shown? 65 6.5 5 1.6

OpenStudy (january123):

https://media.studyisland.com/pics/150680no22.png

OpenStudy (january123):

@Pocarii @Aamirgsen @inowalst

OpenStudy (january123):

@Nnesha @bibby @CloverRacer @iSpiffy2wice

OpenStudy (january123):

@KyanTheDoodle anyone?

OpenStudy (inowalst):

First off, how do you find the mean of a set of data?

OpenStudy (january123):

The mean is the average value of all data in a set. Add up the numbers then divide by the number of values in the set to find the mean.

OpenStudy (january123):

I'm stuck on the dividing part tho

OpenStudy (inowalst):

Alright. So first, lets multiply. \[3 \times 4\]\[4 \times 2\]\[5 \times 2\]\[6 \times 2\]\[7 \times 1\]\[8 \times 2\Find each product, then tell me what they are.

OpenStudy (inowalst):

Last is supposed to be 8 x 2

OpenStudy (january123):

1. 12 2. 8 3. 10 4. 12 5. 7 6. 16

OpenStudy (inowalst):

Alright, excellent. So now, as you told me, add up all the numbers. Then tell me what you get.

OpenStudy (january123):

65 ty :0

OpenStudy (anonymous):

i think 60?

OpenStudy (anonymous):

never mind

OpenStudy (january123):

THAT'S NOT THE ANSWER

OpenStudy (inowalst):

Good, now divide that by 6 and that's your answer.

OpenStudy (january123):

GOOD NIGHT :0 I GTG

OpenStudy (inowalst):

Okay. Goodnight.